OpenStudy (anonymous):

Find all the zeros of the equation. -4x4 - 44x2 + 3600 = 0

OpenStudy (anonymous):

First write the equation as 4x^4 + 44x^2 - 3600 = 0. Next, factor out a 4, you get x^4 + 11x^2 - 900 = 0. Now factor the above.

OpenStudy (john_es):

You can also do the trick, \[z=x^2\Rightarrow z^2+11z-900=0\]And apply, \[x^2=z=\frac{-b\pm\sqrt{b^2+4ac}}{2a}\]of the usual, \[ax^2+bx+c=0\]

OpenStudy (anonymous):

A. 5, -5, 6i, -6i B. 5, 6i C. 5, -5, 6i, 0 D. -5, -6i

OpenStudy (john_es):

The first is the one I would choose.

OpenStudy (anonymous):

z = (-11+-61)/2

OpenStudy (anonymous):

i got c \

OpenStudy (john_es):

Multiply the four roots of A and tell me what number you obtain.

OpenStudy (anonymous):

you can' have just one "i" they come in pairs

OpenStudy (anonymous):

if there is a +i there is a corresponding -i

OpenStudy (john_es):

Remember the trick, the product of the roots must be equal to the independent term (that without x) of the equation.

OpenStudy (anonymous):

you could also try factoring this problem has been cooked up so that \[x^4+11x^2-900\] factors easily

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

25 and 36 can be made to fit nicely